1st Year Course Details 2000-2001

Course  M  Tu W Th F  Start  First Meet
AI1Ah+Bh  4  4  -  4 -  Oct9+Jan29 LTh4/AT + LTh4/AT
Ast1Ah  2  -  -  - 2+4 Oct9  LTh5/AT
Ast1Bh  2+4 -  -  - 2+4 Jan29  LTh1/AT
Ast1Ch  2+4 -  -  - 2  Oct9  LTh3/AT
BusSt1  5  5  -  5 -  Oct9  GeoSqLTh
CL1h  5.10 5.10 - 5.10 -  Oct9  LTh4/AT
CS1Ah  2  11 - 11 -  Oct9  LTh4/AT
CS1Bh  2  11 - 11 -  Jan29  LTh4/AT
EE1h  10  -  - 10 10 Jan29  LTh4/AT
Elect1h  5  5  -  5 -  Jan29  LThB/JCMB
Eng1h  10  10 10 10 10 Oct9  LTh4/AT
HistSci1h  5.10 5.10 - 5.10 -  Oct9  LTh1/AT
HumComm1h  5.10 5.10 - 5.10 -  Jan29  LTh2/AT
IndM1h  9  -  9  - 9  Oct9  LTh1/Sanderson
Logic 1h  5  5  -  5 -  Oct9  LTh3/AT
Ling1  10  10 10 10 10 Oct9  LThB/DHT
Maths Hons 1  varies by module  Oct9  LT2/AT
Maths1Y1h+2h 12  12 12 12 12 Oct9(both)  LTh3/AT
Maths1q  varies by module
Meteor1  10  - 10 - 10 Oct9  LThC/JCMB
PhilSci1h  5  5  -  5 -  Jan29  LTh3/AT3
Physics1Ah+1Bh 11  - 11 - 11 Oct9+Jan29 LTh4/AT + LTh4/AT
Int2Phil1  10  - 10 - 10 Oct9  LThA/DHT
Psych1  11  11 11 11 11 Oct9  LThA/DHT
SciInSoc1Ah  5  5  -  5 -  Jan29  LTh1/AT
StatSoc1h  10  10 - 10 -  Oct9  LTh2/AT
TMSE1h  9  -  9  - 9  Jan29  LTh2/Hudson Beare

2nd Year Course Details 2000-2001

AI2Ah+Bh - 9 - 9 9 Oct10 +Jan30 LTh4AT (first) + D/FH (second)

Business Studies 2 half-courses:

MS&IS  5  5  -  5 -  Oct9  G.08/WRB
OMSC  5  5  -  5 -  Jan29  F.21/Psych
MS&OP  5  5  -  5 -  Jan29  G.19/AFB
CS2  11  - 11 - 11 Oct9+Jan29 LThA/JCMB (both)
EE2  12  12 10 - 10 Oct9  LTh2/Hudson Beare
Ling2  10  10 10 10 10 Oct10  G10/AFB
Maths Hons 2  varies by module  Oct9  LThB/JCMB
Maths2q  varies by module
Maths2Y  12  12 12 12 12 Oct9  LThA/JCMB
Physics2  9  9 9  - 9  Oct9  LThA/JCMB
Phil2Eh  -  2  -  2 2  Jan30  3.01/DHT
Psych 2  11  11 - 11 -  Oct9  G.08/WRB